# MetaWFN: A Platform for Unified Implementation of Many-Electron Wavefunctions

### v1

@@ Abstract
MetaWFN is a C++ template-based architecture designed for flexible and rapid development of wavefunction-based quantum chemical methods. It is highly modular, extendable, and efficient. This is achieved by decoupling the three distinct aspects of quantum chemical methods (i.e., nature of Hamiltonian, structure of wavefunction, and strategy of parallelization ), thereby allowing for separate treatment of them through their internal type-trait and tagging systems furnished by C++ metaprogramming. Once the second-quantized Hamiltonians, whether nonrelativistic (spin-free) or relativistic (spin-dependent), are decomposed into topologically equivalent diagrams for a unified evaluation of the basic coupling coefficients between (randomly selected) spin-free or spin-dependent configuration state functions or Slater determinants incorporating full molecular symmetry (including single or double point group and spin or time reversal symmetry), the many-electron wavefunctions, whether built up with scalar or spinor orbitals, can be assembled with the same templates. As for parallelization, MetaWFN supports both OpenMP and MPI, with the majority of the latter being translated automatically from its OpenMP counterparts. The whole structure of MetaWFN is reviewed here, with some showcases for illustrating its performance.
@@ 1 Introduction
As emphasized recently1, high-precision spectroscopic calculations ought to treat relativistic, correlation, and quantum electrodynamics (QED) effects on an equal footing, so as to match experimental measurements as closely as possible. Since both relativistic and QED effects can be incorporated into second-quantized Hamiltonians2, 3, 4, 5, which take the same form as the nonrelativistic (spin-free) one, it would be nice if all second-quantized Hamiltonians can be manipulated in the same way for subsequent development of wavefunction methods for describing electron correlation. This is indeed possible, thanks to two recent developments: (1) relativistic-QED Hamiltonians, whether four-component or two-component, can be constructed in a unified manner based solely on physical arguments6; (2) all second-quantized Hamiltonians can be decomposed into topologically equivalent diagrams, such that the basic coupling coefficients between (randomly selected) spin-free or spin-dependent configuration state functions (CSF) or Slater determinants (DET) can be evaluated in a unified manner7. What is left is then how to design and implement many-electron wavefunctions for describing electron correlation as accurately as possible. The so-called strongly correlated systems of electrons, such as extended $\pi$ systems and polynuclear transition metal clusters, are particularly intricate in this aspect, due to the fact that the static and dynamic components of the overall correlation are therein strongly entangled and even interchangeable. The currently best methods for such kinds of systems include density matrix renormalization group (DMRG)8, 9, 10, 11 and selected configuration interaction (sCI) plus second-order perturbation theory (sCIPT2)12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 7, 33, 34, 35, 36.
Given so many wavefunction methods, the question is how to implement them in a unified and efficient manner, so as to facilitate the development of new methods. To address this, we first outline the general aspects of many-electron wavefunctions:
@@
    -electron wavefunctions can be expanded in terms of global Hilbert space basis states (HSS) or local Fock space basis states (FSS). Every member of the former carries $N$-electrons and $N$ spin orbitals, whereas that of the latter may have a variable number of electrons and even a variable number of spin orbitals. For instance, the well-established many-body perturbation theory (MBPT), coupled-cluster (CC), and configuration interaction (CI) types of wavefunctions are all built up with an $N$-electron mean-field (Hartree-Fock or CASSCF) reference and additional $N$-electron configurations generated from it by systematic excitations/replacements. No matter how such HSSs are herein truncated, antisymmetry is always maintained automatically. In contrast, when working with FSSs as in DMRG wavefunctions, suitable constraints must be imposed from the outside to conserve the number ($N$) of electrons and antisymmetry. In essence, the two kinds of representations access electron correlation through different routes: the HSS representation tries to improve the initial independent-particle description by selecting low-energy $N$-electron functions from the (first-order) interacting Hilbert space, ending up with a linear combination of limited HSS, whereas the FSS representation attempts to improve the initial independent-subsystem description by selecting high-entanglement local states from the renormalized Fock space, ending up with a linear combination of all FSS. The symmetry adaptation of HSS can be achieved by using appropriate irreducible representations (irrep) of the $U(n)$ and single point groups in the spin-free case, whereas by using time reversal and double point groups in the spin-dependent case. In contrast, the symmetry adaptation of FSS has to be performed in terms of local symmetry operations in both the spin-free37, 38 and spin-dependent39 cases. Additionally, the implementation has to distinguish between complex and real algebras associated with spin-dependent and spin-free wavefunctions, respectively.
@@
    The efficient execution of wavefunction methods relies on proper parallelization and hardware-specific optimizations. Different parallelization patterns (e.g., OpenMP or MPI) or hardware platforms (e.g., CPU or GPU) typically necessitates substantial code rewrites.
To achieve a unified implementation of wavefunction methods, the above aspects must be decoupled as much as possible. To this end, we design here a platform, MetaWFN, which can treat the three aspects separately through their internal type-trait [^1] Type traits are compile-time tools in C++ that provide information about types (e.g., inquiry whether a type represents the non-relativistic Hamiltonian). This enables functions or templates to automatically adjust their behavior based on the types they handle, enhancing flexibility and type safety. and tagging 222 A tag in C++ is an empty struct to label concepts. For example, one can use a struct NonRelativistic to label non-relativistic Hamiltonian. Combined with type-traits, this will enable compilers to process information related to the non-relativistic Hamiltonian. systems furnished by C++ metaprogramming. What is essential is the use of a template to generate temporary subroutines, which are then merged with other relevant codes and compiled. This strategy reduces significantly code duplication and enhances maintainability as the software expands. Use of MetaWFN has been made to achieve a unified implementation7 of the non-relativistic13, 14 and relativistic33, 7 variants of the fully symmetry adapted iCIPT2 (iterative configuration interaction with selection and second-order perturbation) method.
The remaining of the paper is organized as follows. Section sCIPT2 is devoted to general features of the sCIPT2 methods, including unified treatments of Hamiltonians, many-particle basis functions (MPBF), and Hamiltonian matrix elements (HME), selection procedures, and proper implementation of the Epstein-Nesbet second-order perturbation theory (ENPT2)40, 41. The three-layer architecture of MetaWFN is then presented in detail in Section Software Architecture. As a showcase, the computational costs of sf-X2C-iCIPT2, SOiCI, and 4C-iCIPT2 are briefly compared in Section Application. The paper is finally closed with concluding remarks in Section Conclusion and Outlook.
@@
The available nonrelativistic12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 24, 25, 26, 27, 28, 29, 30, 31, 32 and relativistic7, 33, 34, 35, 36 sCIPT2 methods implemented in several program packages42, 43, 44, 45, 46, 47 all involve three major steps, i.e., preparation of a guess space $P_{0}$, iterative update of the variational space until convergence, and second-order perturbation correction to the energy. However, they differ in the actual implementations, which involve the following issues (see Fig. 1):
@@ 2.1 Unified Handling of Hamiltonians
Here, $\{i,j,k,l\}$ refer to molecular orbitals/spinors obtained by a mean-field calculation, whereas $h_{ij}$ and $(ij|kl)$ (in the Mulliken notation) are the integrals of the core Hamiltonian and two-body operators, respectively. It should be emphasized that, in the spin-dependent case, the molecular spinors are assumed here to be symmetrized48 according to both double point group and time reversal symmetries, such that they can be grouped into $n$ Kramers pairs $\{\psi_{i},\psi_{\bar{i}}=\mathcal{K}\psi_{i}\}_{i=1}^{n}$, where $\mathcal{K}$ ($=-i\mathbf{I}_{2}\otimes\sigma_{y}K_{0}$) is the time reversal operator (NB: $K_{0}$ represents complex conjugation). By making only use of the permutation symmetry of the molecular integrals, Eq. (1) can be decomposed as13
@@
$+\sum_{i<j}\left[(ii|jj)E_{ii}E_{ij}+(ij|jj)e_{ij,ji}\right]$ (7)
$H_{2}^{1}$ $=\sum_{i\neq j}\left[(ii|ij)\left(E_{ii}-1\right)E_{ij}+(ij|jj)E_{ij}\left(E_{ij}-1\right)\right]$
@@
$H_{2}^{2}$ $=\sum_{i\leq k}\sum_{j\leq l}^{\prime}\left[2^{-\delta_{k}\delta_{j}}(ij|kl)e_{ij,kl}+\left(1-\delta_{ik}\right)\left(1-\delta_{jl}\right)(il|kj)e_{il,kj}\right]$ (9)
where the superscripts of $H_{i}$ $(i=1,2)$ classify the excitation levels (zero, single or double) between a pair of orbital configurations (CFG), whereas the prime in the summation of $H_{2}^{2}$ indicates that $\{j,l\}\cap\{i,k\}=\varnothing$. The individual terms of $H_{i}^{j}$ ($i=1,2;j=0,1,2$) can be represented pictorially by the diagrams shown in Figs. 2 to 4, which are drawn with the following conventions: (1) The enumeration of orbital/spinor levels starts with zero and increases from bottom to top. (2) The left and right vertices (represented by filled diamonds) indicate creation and annihilation operators, respectively, which form a single generator when connected by a non-vertical line. (3) Products of single generators should always be understood as normal ordered. For instance, Fig. 3(m) means $\{E_{ik}E_{kj}\}=\{E_{kj}E_{ik}\}=E_{ik}E_{kj}-E_{ijq}=e_{ik,kj}$ ($i<j<k$), which is the exchange counterpart of the direct generator $\{E_{ij}E_{kk}\}=\{E_{kk}E_{ij}\}=e_{ij,kkr}$ ($i<j<k$) shown in Fig. 3(g). The same diagrams can be used to compute the basic coupling coefficients (BCC) between spin-free49 or between spin-dependent7 CSFs/DETs. It is just that, in the spin-dependent case, every vertex of the diagrams [i.e., every term in the algebraic expressions (5) to (9)] is composed of both unbarred (A) and barred (B) spinors (e.g., the first term of $H_{2}^{0}$ (7) is composed of 16 terms, in stead of merely one term as in the spin-free case). As a matter of fact, due to the underlying hermiticity, only the $s_{2}$, $c_{x}$ ($x\in[1,6]$), and $d_{x}$ ($x\in[1,7]$) types of diagrams are needed for computing the one- and two-particle BCCs. That is, once the BCCs for these types of diagrams are available, those for the conjugate diagrams can be obtained simply by matrix conjugate transpose. In particular, the tabulated unitary group approach (TUGA)13, 14 can be employed to evaluate and reuse the BCCs between randomly selected spin-free or spin-dependent CSFs/DETs. The decomposed form (4) of the Hamiltonian (1) offers an additional advantage: by using CFGs as the organization unit, the same set of integrals can be utilized for multiple HMEs (vide post). In particular, some upper bounds $\tilde{H}^{IJ}$ for the HMEs between the CSFs/DETs of CFG pairs $|I\rangle$ and $|J\rangle$ can be established14 in advance (cf. Table 1) and then used to facilitate the selection procedure.
It deserves to be mentioned that the above diagrammatic representation of the Hamiltonians can also be used in DMRG working with matrix product states/operators (MPS/MPO). To see this, recall first that the spin-free BCC $\langle I\mu|E_{ij}|J\nu\rangle$ is evaluated in UGA49 via product of the segment values $W_{r}$,
@@
where $\{d_{r}\}_{r=1}^{n}$ are the Shavitt step numbers50 [NB: $b_{r}$ can be derived from $d_{p}$ ($p\in[1,r]$) and $\Delta b_{r}=b_{r}-\tilde{b}_{r}$), in terms of which the bar and ket CSFs can be expressed as $|I\mu\rangle=|\tilde{d}_{1}\cdots\tilde{d}_{n}\rangle$ and $|J\nu\rangle=|d_{1}\cdots d_{n}\rangle$, respectively. In contrast, in the MPO/MPS formalism51 of DMRG, the Hamiltonian and wavefunctions are expressed as
@@
In short, the present diagrammatic representation of the spin-free/dependent Hamiltonians (which include also multi-component Hamiltonians52, 53 and those involving electron-phonon coupling and matter-light interactions54, 55, 56) unifies the evaluation of matrix elements via Slater-Condon rules, UGA, and MPO/MPS. However, the formulation of spin-adapted MPOs/MPSs in terms of the Shavitt graphs50 remains to be further investigated.
$i$$i$ $i$$i$ $i$$i$$j$$j$ $i$$i$$j$$j$
(a) w1 (b) w2 (c) w3 (d) b7=c7
Figure 2: Diagrammatic representation of $H_{1}^{0}$ (a) and $H_{2}^{0}$ (b-d)
$i$$j$ $j$$i$ $i$$i$$j$ $j$$i$$i$ $i$$j$$j$ $j$$i$$j$
(a) s1 (b) s2 (c) s3 (d) s4 (e) s5 (f) s6
$k$$j$$i$$k$ $k$$j$$i$$k$ $i$$k$$k$$j$ $i$$k$$k$$j$ $i$$k$$k$$j$ $i$$k$$k$$j$
(g) s7 (h) s8 (i) s9 (j) s10 (k) s11 (l) s12
$k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$
(m) b6 (n) c6 (o) b4 (p) c4 (q) a2 (r) d2
Figure 3: Diagrammatic representation of $H_{1}^{1}$ (a-b) and $H_{2}^{1}$ (c-r)
$i$$k$$j$$l$ $i$$k$$j$$l$ $j$$l$$i$ $i$$k$$j$$l$ $i$$j$$k$ $j$$i$
(a) a1 (b) a3 (c) a4 (d) a5 (e) a6 (f) a7
$i$$k$$j$$l$ $j$$l$$i$ $i$$k$$j$$l$ $i$$k$$j$$l$
(g) b1 (h) b2 (i) b3 (j) b5
$i$$k$$j$$l$ $i$$k$$j$ $i$$k$$j$$l$ $i$$k$$j$$l$
(k) c1 (l) c2 (m) c3 (n) c5
$i$$k$$j$$l$ $i$$k$$j$$l$ $i$$k$$j$ $i$$k$$j$$l$ $j$$l$$i$ $j$$i$
(o) d1 (p) d3 (q) d4 (r) d5 (s) d6 (t) d7
Figure 4: Diagrammatic representation of $H_{2}^{2}$
Table 1: Correspondence between UGA Diagrams (see Figs. 2 to 4) and integral categories (IntC) as well as Upper Bounds $\tilde{H}^{IJ}$ for Hamiltonian Matrix Elements over CSFs/DETs

### v2

# MetaWave: A Platform for Unified Implementation of Nonrelativistic and Relativistic Wavefunctions

@@ Abstract
MetaWave is a C++ template-based architecture designed for unified implementation of nonrelativistic and relativistic wavefunction-based quantum chemical methods. It is highly modular, extendable, and efficient. This is achieved by decoupling the three distinct aspects of quantum chemical methods (i.e., nature of Hamiltonian, structure of wavefunction, and strategy of parallelization ), thereby allowing for separate treatment of them through their internal type-trait and tagging systems furnished by C++ metaprogramming. Once the second-quantized Hamiltonians, whether nonrelativistic (spin-free) or relativistic (spin-dependent), are decomposed into topologically equivalent diagrams for a unified evaluation of the basic coupling coefficients between (randomly selected) spin-free or spin-dependent configuration state functions or Slater determinants incorporating full molecular symmetry (including single or double point group and spin or time reversal symmetry), the many-electron wavefunctions, whether built up with scalar or spinor orbitals, can be assembled with the same templates. As for parallelization, MetaWave supports both OpenMP and MPI, with the majority of the latter being translated automatically from its OpenMP counterparts. The whole structure of MetaWave is reviewed here, with some showcases for illustrating its performance.
@@ 1 Introduction
As emphasized recently1, high-precision spectroscopic calculations ought to treat relativistic, correlation, and quantum electrodynamics (QED) effects on an equal footing, so as to match experimental measurements as closely as possible. Since both relativistic and QED effects can be incorporated into second-quantized Hamiltonians2, 3, 4, 5, which take the same form as the nonrelativistic (spin-free) one, it would be nice if all second-quantized Hamiltonians can be manipulated in the same way for subsequent development of wavefunction methods for describing electron correlation. This is indeed possible, thanks to two recent developments: (1) relativistic-QED Hamiltonians, whether four-component or two-component, can be constructed in a unified manner based solely on physical arguments6; (2) all second-quantized Hamiltonians can be decomposed into topologically equivalent diagrams, such that the basic coupling coefficients between (randomly selected) spin-free or spin-dependent configuration state functions (CSF) or Slater determinants (DET) can be evaluated in a unified manner7. What is left is then how to design and implement many-electron wavefunctions for describing electron correlation as accurately as possible. The so-called strongly correlated systems of electrons, such as extended $\pi$ systems and polynuclear transition metal clusters, are particularly intricate in this aspect, due to the fact that the static and dynamic components of the overall correlation are therein strongly entangled and even interchangeable. The currently best methods for such kinds of systems include density matrix renormalization group (DMRG)8, 9, 10, 11, 12, 13, 14 and selected configuration interaction (sCI) plus second-order perturbation theory (sCIPT2)15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 7, 36, 37, 38, 39.
Given so many wavefunction methods, the question is how to extend them to the relativistic realm, yet without significant code rewriting. To address this, we first outline the general aspects of wavefunctions methods:
@@
    -electron wavefunctions can be expanded in terms of Hilbert space basis states (HSS) or Fock space basis states (FSS). Every member of the former carries $N$-electrons and $N$ spin orbitals, whereas that of the latter may have a variable number of electrons and even a variable number of spin orbitals. For instance, the well-established many-body perturbation theory (MBPT), coupled-cluster (CC), and configuration interaction (CI) types of wavefunctions are all built up with an $N$-electron mean-field (Hartree-Fock or CASSCF) reference and additional $N$-electron configurations generated from it by systematic excitations/replacements. No matter how such HSSs are herein truncated, antisymmetry is always maintained automatically. In contrast, when working with local FSSs as in DMRG wavefunctions, suitable constraints must be imposed from the outside to conserve the number ($N$) of electrons and antisymmetry. In essence, the two kinds of representations access electron correlation through different routes: the HSS representation tries to improve the initial independent-particle description by selecting low-energy $N$-electron functions from the (first-order) interacting Hilbert space, ending up with a linear (in CI) or nonlinear (in CC) combination of limited HSS, whereas the local FSS representation attempts to improve the initial independent-subsystem description by selecting high-entanglement local states from the renormalized Fock space, ending up with a particular linear combination of all HSS. The symmetry adaptation of HSS can be achieved by using appropriate irreducible representations (irrep) of the $U(n)$ and single point groups in the spin-free case, whereas by using time reversal and double point groups in the spin-dependent case. In contrast, the symmetry adaptation of local FSS has to be performed in terms of local symmetry operations in both the spin-free40, 41 and spin-dependent42 cases. Additionally, the implementation has to distinguish between complex and real algebras associated with spin-dependent and spin-free wavefunctions, respectively.
@@
    The efficient execution of wavefunction methods relies on proper parallelization and hardware-specific optimizations. Different parallelization patterns (e.g., OpenMP or MPI) or hardware platforms (e.g., CPU, GPU or their hybrid43, 44) typically necessitates substantial code rewrites.
To achieve a unified implementation of wavefunction methods, the above aspects must be decoupled as much as possible. To this end, we design here a platform, MetaWave, which can treat the three aspects separately through their internal type-trait [^1] Type traits are compile-time tools in C++ that provide information about types (e.g., inquiry whether a type represents the nonrelativistic Hamiltonian). This enables functions or templates to automatically adjust their behavior based on the types they handle, enhancing flexibility and type safety. and tagging 222 A tag in C++ is an empty struct to label concepts. For example, one can use a struct NonRelativistic to label nonrelativistic Hamiltonian. Combined with type-traits, this will enable compilers to process information related to the nonrelativistic Hamiltonian. systems furnished by C++ metaprogramming. What is essential is the use of a template to generate temporary subroutines, which are then merged with other relevant codes and compiled. This strategy reduces significantly code duplication and enhances maintainability as the software expands. Use of MetaWave has been made to achieve a unified implementation7 of the nonrelativistic16, 17 and relativistic36, 7 variants of the fully symmetry adapted iCIPT2 (iterative configuration interaction with selection and second-order perturbation) method. Note in passing that the idea of decoupling different aspects of wavefunction methods was already adopted in the HORTON program package45, where the Fanpy library46 is also highly modular and general for implementing new wavefunction methods based on the flexible ansätze for $N$-electron configuration interaction (FANCI)47 and perturbation theory (FANPT) frameworks48. However, Fanpy is a research tool written purely in Python, with the primary goal being quick implementation and test of a new wavefunction ansatz, instead of high-performance parallel computing. Moreover, Fanpy does not support relativistic Hamiltonians and wavefunctions, which are the primary concern of MetaWave.
The remaining of the paper is organized as follows. Sec. 2 is devoted to general features of the sCIPT2 methods, including unified treatments of Hamiltonians, many-particle basis functions (MPBF), and Hamiltonian matrix elements (HME), selection procedures, and proper implementation of the Epstein-Nesbet second-order perturbation theory (ENPT2)49, 50. The three-layer architecture of MetaWave is then presented in detail in Sec. 3. As a showcase, the computational costs of sf-X2C-iCIPT2, SOiCI, and 4C-iCIPT2 are briefly compared in Sec. 4. The paper is finally closed with concluding remarks in Sec. 5.
@@
The available nonrelativistic15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 27, 28, 29, 30, 31, 32, 33, 34, 35 and relativistic7, 36, 37, 38, 39 sCIPT2 methods implemented in several program packages51, 52, 53, 54, 55, 56 all involve three major steps, i.e., preparation of a guess space $P_{0}$, iterative update of the variational space until convergence, and second-order perturbation correction to the energy. However, they differ in the actual implementations, which involve the following issues (see Fig. 1):
@@ 2.1 Unified Handling of Hamiltonians
Here, $\{i,j,k,l\}$ refer to molecular orbitals/spinors obtained by a mean-field calculation, whereas $h_{ij}$ and $(ij|kl)$ (in the Mulliken notation) are the integrals of the core Hamiltonian and two-body operators, respectively. It should be emphasized that, in the spin-dependent case, the molecular spinors are assumed here to be symmetrized57 according to both double point group and time reversal symmetries, such that they can be grouped into $n$ Kramers pairs $\{\psi_{i},\psi_{\bar{i}}=\mathcal{K}\psi_{i}\}_{i=1}^{n}$, where $\mathcal{K}$ ($=-i\mathbf{I}_{2}\otimes\sigma_{y}K_{0}$) is the time reversal operator (NB: $K_{0}$ represents complex conjugation). By making only use of the permutation symmetry of the molecular integrals, Eq. (1) can be decomposed as16
@@
$+\sum_{i<j}\left[(ii|jj)E_{ii}E_{jj}+(ij|jj)e_{ij,ji}\right]$ (7)
$H_{2}^{1}$ $=\sum_{i\neq j}\left[(ii|ij)\left(E_{ii}-1\right)E_{ij}+(ij|jj)E_{ij}\left(E_{jj}-1\right)\right]$
@@
$H_{2}^{2}$ $=\sum_{i\leq k}\sum_{j\leq l}^{\prime}\left[2^{-\delta_{ik}\delta_{jl}}(ij|kl)e_{ij,kl}+\left(1-\delta_{ik}\right)\left(1-\delta_{jl}\right)(il|kj)e_{il,kj}\right]$ (9)
where the superscripts of $H_{i}$ $(i=1,2)$ classify the excitation levels (zero, single or double) between a pair of orbital configurations (oCFG), whereas the prime in the summation of $H_{2}^{2}$ indicates that $\{j,l\}\cap\{i,k\}=\varnothing$. The individual terms of $H_{i}^{j}$ ($i=1,2;j=0,1,2$) can be represented pictorially by the diagrams shown in Figs. 2 to 4, which are drawn with the following conventions: (1) The enumeration of orbital/spinor levels starts with zero and increases from bottom to top. (2) The left and right vertices (represented by filled diamonds) indicate creation and annihilation operators, respectively, which form a single generator when connected by a non-vertical line. (3) Products of single generators should always be understood as normal ordered. For instance, Fig. 3(m) means $\{E_{ik}E_{kj}\}=\{E_{kj}E_{ik}\}=E_{ik}E_{kj}-E_{ij}=e_{ik,kj}$ ($i<j<k$), which is the exchange counterpart of the direct generator $\{E_{ij}E_{kk}\}=\{E_{kk}E_{ij}\}=e_{ij,kk}$ ($i<j<k$) shown in Fig. 3(g). The same diagrams can be used to compute the basic coupling coefficients (BCC) between spin-free58 or between spin-dependent7 CSFs/DETs. It is just that, in the spin-dependent case, every vertex of the diagrams [i.e., every term in the algebraic expressions (5) to (9)] is composed of both unbarred (A) and barred (B) spinors (e.g., the first term of $H_{2}^{0}$ (7) is composed of 16 terms, in stead of merely one term as in the spin-free case). As a matter of fact, due to the underlying hermiticity, only the $s_{2}$, $c_{x}$ ($x\in[1,6]$), and $d_{x}$ ($x\in[1,7]$) types of diagrams are needed for computing the one- and two-particle BCCs. That is, once the BCCs for these types of diagrams are available, those for the conjugate diagrams can be obtained simply by matrix conjugate transpose. In particular, the tabulated unitary group approach (TUGA)16, 17 can be employed to evaluate and reuse the BCCs between randomly selected spin-free or spin-dependent CSFs/DETs. The decomposed form (4) of the Hamiltonian (1) offers an additional advantage: by using oCFGs as the organization unit, the same set of integrals can be utilized for multiple HMEs (vide post). In particular, some upper bounds $\tilde{H}^{IJ}$ for the HMEs between the CSFs/DETs of oCFG pairs $|I\rangle$ and $|J\rangle$ can be established17 in advance (cf. Table 1) and then used to facilitate the selection procedure.
It deserves to be mentioned that the above diagrammatic representation of the Hamiltonians can also be used in DMRG working with matrix product states/operators (MPS/MPO). To see this, recall first that the spin-free BCC $\langle I\mu|E_{ij}|J\nu\rangle$ is evaluated in UGA58 via product of the segment values $W_{r}$,
@@
where $\{d_{r}\}_{r=1}^{n}$ are the Shavitt step numbers59 [NB: $b_{r}$ can be derived from $d_{p}$ ($p\in[1,r]$) and $\Delta b_{r}=b_{r}-\tilde{b}_{r}$), in terms of which the bar and ket CSFs can be expressed as $|I\mu\rangle=|\tilde{d}_{1}\cdots\tilde{d}_{n}\rangle$ and $|J\nu\rangle=|d_{1}\cdots d_{n}\rangle$, respectively. In contrast, in the MPO/MPS formalism60 of DMRG, the Hamiltonian and wavefunctions are expressed as
@@
In short, the present diagrammatic representation of the spin-free/dependent Hamiltonians (which can be extended to include multicomponent Hamiltonians61, 62 and those involving electron-phonon coupling and matter-light interactions63, 64, 65) unifies the evaluation of matrix elements via Slater-Condon rules, UGA, and MPO/MPS. Nevertheless, the formulation of spin-adapted MPOs/MPSs in terms of the Shavitt graphs59 remains to be further investigated. Moreover, it should be mentioned that some tensor contraction engine should further be incorporated into MetaWave for unified implementation of relativistic and nonrelativistic CC methods.
Table 1: Correspondence between UGA Diagrams58 (see Figs. 2 to 4) and integral categories (IntC) as well as Upper Bounds $\tilde{H}^{IJ}$ for Hamiltonian Matrix Elements over CSFs/DETs